OpenStudy (anonymous):

Quick question: when using sigma notation, when you shift the lower index by x amount do you also have to add or subtract that amount from the upper index?

OpenStudy (0bucchr):

nope but you do have to alter to values inside in sigma

OpenStudy (anonymous):

\[\sum_{k=1}^{n} = ((n-1)!/(k-1)!(n-k-1)!)*p^{k-1}(1-p)^{n-k}\] There we go, if I wanted to shift to k=0 how would it change?

OpenStudy (amistre64):

lets see what we do to a simple example when n=3, 2+n = 5 say we shift the index (n) by -1 but we want to retain the same value when n=3-1, 2+(n+1) = 5

OpenStudy (0bucchr):

basically decrease index increase value by same amount inside usually you can tell it will be good. Look at all those -1's falling away

OpenStudy (amistre64):

if we desire to shift the index by "a"; but retain the same value then n= 3+a; 2+(n-a) = 5

OpenStudy (anonymous):

I need to get rid of the n-1's as well, I was hoping simply changing the lower index would do that

OpenStudy (anonymous):

That is, they need to be just n's

OpenStudy (amistre64):

only your "k" parts are affected; since thats the driving part of the summation

OpenStudy (anonymous):

How would I affect the n parts? substitution?

OpenStudy (amistre64):

this seems to be part of a larger question; what is the bigger question that you are trying to work out?

OpenStudy (0bucchr):

but you can change (n-1)! into n!/n-1

OpenStudy (amistre64):

the "n"s are just the stopping value; you cant affect them that I can recall

OpenStudy (0bucchr):

in the same way that you can change (n+1)! into (n+1)!*n!

OpenStudy (amistre64):

\[\sum_{k=1}^{n} = \frac{(n-1)!}{(k-1)!!(n-k-1)!}~p^{k-1}(1-p)^{n-k}\] \[\sum_{k=1-1}^{n} = \frac{(n-1)!}{(k+1-1)!(n-(k+1)-1)!}~p^{k+1-1}(1-p)^{n-(k+1)}\] \[\sum_{k=0}^{n} = \frac{(n-1)!}{(k)!(n-k-2))!}~p^{k}(1-p)^{n-k-1}\] hmmm

OpenStudy (amistre64):

if the upper bound was "fixed", then you could adjust it; but as is, there is no specific stopping point on it to adjust

OpenStudy (anonymous):

I made a mistake \[\sum_{k=1}^{n} = ((n-1)!/(k-1)!(n-k)!)*p^{k-1}(1-p)^{n-k}\] thats what it should be

OpenStudy (amistre64):

\[\sum_{k=1}^{n} = \frac{(n-1)!}{(k-1)!(n-k)!}~p^{k-1}(1-p)^{n-k}\] \[\sum_{k=0}^{n} = \frac{(n-1)!}{(k)!(n-k-1)!}~p^{k}(1-p)^{n-k-1}\] still not working out the way you want it too :/

OpenStudy (anonymous):

and trying to get to \[\sum_{k=0}^{n} = (n!/k!(n-k)!)*p^{k}(1-p)^{n-k}\]

OpenStudy (anonymous):

which is the binomial distribution

OpenStudy (amistre64):

you might have made an error getting it to this point. This is not the start of the question is it???

OpenStudy (anonymous):

Im trying to prove E(X) = np for binomial but Ive gotten to this point, so close, just need to show that sum is \[\sum_{k=0}^{n} = (n!/k!(n-k)!)*p^{k}(1-p)^{n-k}\] which is 1 and then Id be done

OpenStudy (anonymous):

not sure how to manipulate it further

OpenStudy (amistre64):

http://www.math.ubc.ca/~feldman/m302/binomial.pdf this might be a better read for you then

OpenStudy (anonymous):

ah I did need substitution, thanks for that link, Ill see if I can keep going knowing what substitution to make.